# Lie algebras generated by reflections in types BCD

### v1

@@ 4.2. Lie algebras generated by reflections
The structure of $\mathfrak{s}_{n}$ was determined by Marin [13]. Let $\mathcal{X}_{n}=\sum_{j=1}^{n}X_{j}\in\mathfrak{d}_{n}\subseteq\mathfrak{b}_{n}$ be the sum of the YJM elements, and let $\mathcal{T}_{n}=\sum_{j=1}^{n}t_{j}\in\mathfrak{b}_{n}$. Then Lemmas 4.1.2 and 4.1.4 give:
@@
The next proposition extends [13, Proposition 2] to types B and D.
@@
Let $n\geq 2$, let $(G,\mathfrak{g})$ be either the pair $(\mathcal{B}_{n},\mathfrak{b}_{n})$ or $(\mathcal{D}_{n},\mathfrak{d}_{n})$, and let $V_{1}$ and $V_{2}$ be two simple $\mathbb{C}G$-modules of dimension greater than $1$. [^1] All simple $\mathbb{C}\mathcal{D}_{2}$-modules are one-dimensional, so for $(G,\mathfrak{g})=(\mathcal{D}_{n},\mathfrak{d}_{n})$ this implies that $n\geq 3$. Consider the following statements:
@@
Statements (1) and (2) are equivalent because $\mathfrak{g}$ contains a set of associative algebra generators for $\mathbb{C}G$, and (2) evidently implies (3). Now suppose (3) holds. Let $\phi:V_{1}\to V_{2}$ be a $\mathfrak{g}^{\prime}$-module isomorphism, and let $\rho_{1}:\mathbb{C}G\to\operatorname{End}(V_{1})$ and $\rho_{2}:\mathbb{C}G\to\operatorname{End}(V_{2})$ be the $\mathbb{C}G$-module structure maps for $V_{1}$ and $V_{2}$, respectively. Then for all $x\in\mathfrak{g}^{\prime}$, one has $\phi\circ\rho_{1}(x)=\rho_{2}(x)\circ\phi$. Let $C_{(21^{n-2},\emptyset)}$ be the first conjugacy class identified in (2.1.1). The elements of $C_{(21^{n-2},\emptyset)}$ generate the subgroup $\mathcal{D}_{n}$ of $G$ (which may be equal to $G$). Then one can argue as in the proof of [13, Proposition 2] (cf. also the proof of [3, Proposition 4.6.3]) to show that either:
@@
Now since (Iso1) holds, it follow by multiplicativity that $\phi\circ\rho_{1}(g)=\rho_{2}(g)\circ\phi$ for all $g\in\mathcal{D}_{n}$. Then the proposition is true if $G=\mathcal{D}_{n}$, so suppose that $G=\mathcal{B}_{n}$. Let $C_{(1^{n-1},1)}$ be the second conjugacy class identified in (2.1.1). Then one can again argue as in the proof of [13, Proposition 2] to show that either:
@@
If (Iso3) holds, then one deduces by multiplicativity that $\phi\circ\rho_{1}(g)=\rho_{2}(g)\circ\phi$ for all $g\in\mathcal{B}_{n}$, and hence $V_{1}\cong V_{2}$ as $\mathbb{C}\mathcal{B}_{n}$-modules. So suppose (Iso4) holds. In this case the scalars $b_{1},b_{2}$ must each be either $+1$ or $-1$, since each simple $\mathbb{C}\mathcal{B}_{n}$-module decomposes into ±$\pm 1$-eigenspaces for the action of the elements $t\in C_{(1^{n-1},1)}$. Since each $t\in C_{(1^{n-1},1)}$ is acting on $V_{1}$ by the same scalar value, we get from the descriptions in Sections 2.2 and 2.3 that $V_{1}\cong S^{(\lambda,\emptyset)}$ or $V_{1}\cong S^{(\emptyset,\lambda)}$ for some partition $\lambda\vdash n$, and similarly $V_{2}\cong S^{(\mu,\emptyset)}$ or $V_{2}\cong S^{(\emptyset,\mu)}$ for some $\mu\vdash n$. Condition (Iso1) implies that $V_{1}\cong V_{2}$ as $\mathbb{C}\mathfrak{S}_{n}$-modules, so $\lambda=\mu$. Thus either $V_{1}\cong V_{2}$, or $\{V_{1},V_{2}\}\cong\{S^{(\lambda,\emptyset)},S^{(\emptyset,\lambda)}\}$.
@@ 4.3.1. Dual pairs
Let $(\lambda,\mu)\in\mathcal{BP}(n)$. As remarked after Theorem 2.2.2, all simple $\mathbb{C}\mathcal{B}_{n}$-modules are self-dual. Then from (2.2.5) and Lemma 4.1.6, one gets $[S^{(\lambda,\mu)}]^{*,\operatorname{Lie}}\cong S^{(\mu^{*},\lambda^{*})}$ as $\mathfrak{b}_{n}$-modules. Thus, with respect to an appropriate choice of basis, the map $\rho_{(\mu^{*},\lambda^{*})}:\mathfrak{b}_{n}\to\mathfrak{gl}(\mu^{*},\lambda^{*})$ can be written in the form $X\mapsto-\rho_{(\lambda,\mu)}(X)^{t}$, where $A^{t}$ denotes the transpose of $A$. This map evidently factors through $\rho_{(\lambda,\mu)}$, and one gets the first commutative diagram in (4.3.3) below. Using Remark 3.2.3, one similarly sees that $[S^{\left\{\lambda,\mu\right\}}]^{*,\operatorname{Lie}}\cong S^{\left\{\lambda^{*},\mu^{*}\right\}}$ as $\mathfrak{d}_{n}$-modules, and deduces that $\rho_{\left\{\lambda^{*},\mu^{*}\right\}}$ factors through $\rho_{\left\{\lambda,\mu\right\}}$, yielding the second commutative diagram in (4.3.3).
@@
$[S^{\left\{\lambda,\pm\right\}}]^{*,\operatorname{Lie}}\cong\begin{cases}S^{\left\{\lambda^{*},\pm\right\}}&\text{if $n/2$ is even,}\\ S^{\left\{\lambda^{*},\mp\right\}}&\text{if $n/2$ is odd.}\end{cases}$
@@ 4.3.2. Self-dual Lie algebra representations
The arguments in this section are based on those in [11, III.3.2.3]. First let $(\lambda,\mu)\in\mathcal{BP}(n)$ with $(\lambda,\mu)=(\mu^{*},\lambda^{*})$. Then $\mu=\lambda^{*}$, $n$ is even, and $\lambda$ and $\mu$ are partitions of $n/2$. Let $\langle-,-\rangle_{(\lambda,\mu)}$ be the bilinear form on $S^{(\lambda,\mu)}$ discussed in Remark 2.4.2, and let $(-|-)_{(\lambda,\mu)}$ be the nondegenerate bilinear form on $S^{(\lambda,\mu)}$ defined in terms of the linear isomorphism $\phi_{\varepsilon}=\phi_{\varepsilon}^{(\lambda,\mu)}:S^{(\lambda,\mu)}\to S^{(\lambda,\mu)}$ of Remark 2.4.5 by
@@
The bilinear form (|)(-|-)( - | - ) is symmetric if $n/2$ is even and is anti-symmetric if $n/2$ is odd. Thus,
@@
[…]
@@
Since $\varepsilon(\sigma)=(-1)^{n/2}$, this implies that the form (|)(-|-)( - | - ) is symmetric if $n/2$ is even, and is anti-symmetric if $n/2$ is odd. Finally, (|)(-|-)( - | - ) is non-degenerate because $\langle$-,$\rangle$ - , - ⟩ is non-degenerate and $\phi_{\varepsilon}$ is a linear isomorphism. ∎
Next let $\left\{\lambda,\mu\right\}\in\mathcal{BP}\{n\}$ with $\left\{\lambda,\mu\right\}=\left\{\lambda^{*},\mu^{*}\right\}$. If $(\lambda,\mu)=(\mu^{*},\lambda^{*})$, then $\mathfrak{d}_{\left\{\lambda,\mu\right\}}\subseteq\mathfrak{osp}\{\lambda,\mu\}$ as defined in Lemma 4.3.1, so suppose that $(\lambda,\mu)=(\lambda^{*},\mu^{*})$. Let (|){$(-|-)_{\left\{\lambda,\mu\right\}}$ be the nondegenerate bilinear form on $S^{\left\{\lambda,\mu\right\}}=\operatorname{Res}^{\mathcal{B}_{n}}_{\mathcal{D}_{n}}(S^{(\lambda,\mu)})$ defined via the map $\phi_{\varepsilon^{\prime\prime}}=\phi_{\varepsilon^{\prime\prime}}^{(\lambda,\mu)}$ of Lemma 2.4.4 by
@@
Let $(\lambda,\mu)\in\mathcal{BP}(n)$ with $(\lambda,\mu)=(\lambda^{*},\mu^{*})$, and let (|)=(|){$(-|-)=(-|-)_{\left\{\lambda,\mu\right\}}$ be the bilinear form on $S^{\left\{\lambda,\mu\right\}}$ defined by (4.3.9). Then
@@
Let $b(\lambda)$ and $b(\mu)$ be the lengths of the diagonals of the Young diagrams of $\lambda$ and $\mu$, respectively. Then (|)(-|-)( - | - ) is symmetric if $\frac{n-b(\lambda)-b(\mu)}{2}$ is even, and is anti-symmetric otherwise. [^2] If $\lambda=\lambda^{*}$ and $\mu=\mu^{*}$, then $|\lambda|-b(\lambda)$ and $|\mu|-b(\mu)$ must each be even, hence their sum $n-b(\lambda,\mu)$ is even. Thus
@@
The proof is parallel to that of Lemma 4.3.1. In determining whether (|)(-|-)( - | - ) is symmetric or anti-symmetric, let $\sigma\in\mathfrak{S}_{n}\subset\mathcal{D}_{n}$ be the permutation that maps $T\in\mathbb{T}(\lambda,\mu)$ to $T^{*}$. Then $\sigma$ is a product of $\frac{|\lambda|-b(\lambda)}{2}+\frac{|\mu|-b(\mu)}{2}$ disjoint transpositions, independent of $T$, so $\varepsilon^{\prime\prime}(\sigma)=1$ if $\frac{n-b(\lambda)-b(\mu)}{2}$ is even, and $\varepsilon^{\prime\prime}(\sigma)=-1$ if $\frac{n-b(\lambda)-b(\mu)}{2}$ is odd. ∎
@@
Let $\left\{\lambda,\lambda\right\}\in\mathcal{BP}\{n\}$ with $\lambda=\lambda^{*}$ and $n/2$ even, and let (|)=(|){$(-|-)=(-|-)_{\left\{\lambda,\pm\right\}}$ be the restriction to $S^{\left\{\lambda,\pm\right\}}$ of the bilinear form on $S^{\left\{\lambda,\lambda\right\}}$ defined by (4.3.9). Then
@@
Let $b(\lambda)$ be the length of the diagonal of the Young diagram of $\lambda$. Then (|)(-|-)( - | - ) is symmetric if $b(\lambda)$ is even, and is anti-symmetric if $b(\lambda)$ is odd. Thus
@@
If $\left\{\lambda,\lambda\right\}\in\mathcal{BP}\{n\}$ and $\lambda=\lambda^{*}$, then (4.3.7) and (4.3.9) define two different bilinear forms on $S^{\left\{\lambda,\lambda\right\}}=\operatorname{Res}^{\mathcal{B}_{n}}_{\mathcal{D}_{n}}(S^{(\lambda,\lambda)})$. For $n/2$ even, one can check that with respect to either form, the subspaces $S^{\left\{\lambda,+\right\}}$ and $S^{\left\{\lambda,-\right\}}$ of $S^{\left\{\lambda,\lambda\right\}}$ are orthogonal, the restrictions of the two forms to $S^{\left\{\lambda,+\right\}}$ are the same, and the restrictions of the two forms to $S^{\left\{\lambda,-\right\}}$ differ by the scalar $-1$. Thus, for $n/2$ even, the ambiguity in definition does not affect the conclusions of Lemma 4.3.3. For $n/2$ odd, one gets that with respect to either form, the spaces $S^{\left\{\lambda,+\right\}}$ and $S^{\left\{\lambda,-\right\}}$ are each totally isotropic, and the forms each define perfect pairings (differing by the scalar $-1$) between $S^{\left\{\lambda,+\right\}}$ and $S^{\left\{\lambda,-\right\}}$.

### v2

@@ 4.2. Lie algebras generated by reflections
The structure of $\mathfrak{s}_{n}$ was determined by Marin [Marin:2007]. Let $\mathcal{X}_{n}=\sum_{j=1}^{n}X_{j}\in\mathfrak{d}_{n}\subseteq\mathfrak{b}_{n}$ be the sum of the YJM elements, and let $\mathcal{T}_{n}=\sum_{j=1}^{n}t_{j}\in\mathfrak{b}_{n}$. Then Lemmas 4.1.2 and 4.1.4 give:
@@
The next proposition extends [Marin:2007, Proposition 2] to types B and D.
@@
Let $n\geq 2$, let $(G,\mathfrak{g})$ be either the pair $(\mathcal{B}_{n},\mathfrak{b}_{n})$ or $(\mathcal{D}_{n},\mathfrak{d}_{n})$, and let $V_{1}$ and $V_{2}$ be two simple $\mathbb{C}G$-modules of dimension greater than 11.111All simple $\mathbb{C}\mathcal{D}_{2}$-modules are one-dimensional, so for $(G,\mathfrak{g})=(\mathcal{D}_{n},\mathfrak{d}_{n})$ this implies that $n\geq 3$. Consider the following statements:
@@
Statements (1) and (2) are equivalent because $\mathfrak{g}$ contains a set of associative algebra generators for $\mathbb{C}G$, and (2) evidently implies (3). Now suppose (3) holds. Let $\phi:V_{1}\to V_{2}$ be a $\mathfrak{g}^{\prime}$-module isomorphism, and let $\rho_{1}:\mathbb{C}G\to\operatorname{End}(V_{1})$ and $\rho_{2}:\mathbb{C}G\to\operatorname{End}(V_{2})$ be the $\mathbb{C}G$-module structure maps for $V_{1}$ and $V_{2}$, respectively. Then for all $x\in\mathfrak{g}^{\prime}$, one has $\phi\circ\rho_{1}(x)=\rho_{2}(x)\circ\phi$. Let $C_{(21^{n-2},\emptyset)}$ be the first conjugacy class identified in (2.1.1). The elements of $C_{(21^{n-2},\emptyset)}$ generate the subgroup $\mathcal{D}_{n}$ of $G$ (which may be equal to $G$). Then one can argue as in the proof of [Marin:2007, Proposition 2] (cf. also the proof of [DK:2023, Proposition 4.6.3]) to show that either:
@@
Now since (Iso1) holds, it follow by multiplicativity that $\phi\circ\rho_{1}(g)=\rho_{2}(g)\circ\phi$ for all $g\in\mathcal{D}_{n}$. Then the proposition is true if $G=\mathcal{D}_{n}$, so suppose that $G=\mathcal{B}_{n}$. Let $C_{(1^{n-1},1)}$ be the second conjugacy class identified in (2.1.1). Then one can again argue as in the proof of [Marin:2007, Proposition 2] to show that either:
@@
If (Iso3) holds, then one deduces by multiplicativity that $\phi\circ\rho_{1}(g)=\rho_{2}(g)\circ\phi$ for all $g\in\mathcal{B}_{n}$, and hence $V_{1}\cong V_{2}$ as $\mathbb{C}\mathcal{B}_{n}$-modules. So suppose (Iso4) holds. In this case the scalars $b_{1},b_{2}$ must each be either +1+1 or 1-1, since each simple $\mathbb{C}\mathcal{B}_{n}$-module decomposes into $\pm 1$-eigenspaces for the action of the elements $t\in C_{(1^{n-1},1)}$. Since each $t\in C_{(1^{n-1},1)}$ is acting on $V_{1}$ by the same scalar value, we get from the descriptions in Sections 2.2 and 2.3 that $V_{1}\cong S^{(\lambda,\emptyset)}$ or $V_{1}\cong S^{(\emptyset,\lambda)}$ for some partition $\lambda\vdash n$, and similarly $V_{2}\cong S^{(\mu,\emptyset)}$ or $V_{2}\cong S^{(\emptyset,\mu)}$ for some $\mu\vdash n$. Condition (Iso1) implies that $V_{1}\cong V_{2}$ as $\mathbb{C}\mathfrak{S}_{n}$-modules, so $\lambda=\mu$. Thus either $V_{1}\cong V_{2}$, or $\{V_{1},V_{2}\}\cong\{S^{(\lambda,\emptyset)},S^{(\emptyset,\lambda)}\}$.
@@ 4.3.1. Dual pairs
Let $(\lambda,\mu)\in\mathcal{BP}(n)$. As remarked after Theorem 2.2.2, all simple $\mathbb{C}\mathcal{B}_{n}$-modules are self-dual. Then from (2.2.5) and Lemma 4.1.6, one gets $S^{(\lambda,\mu)}]^{*,\operatorname{Lie}}\cong S^{(\mu^{*},\lambda^{*})}$ as $\mathfrak{b}_{n}$-modules. Thus, with respect to an appropriate choice of basis, the map $\rho_{(\mu^{*},\lambda^{*})}:\mathfrak{b}_{n}\to\mathfrak{gl}(\mu^{*},\lambda^{*})$ can be written in the form $X\mapsto-\rho_{(\lambda,\mu)}(X)^{t}$, where $A^{t}$ denotes the transpose of $A$. This map evidently factors through $\rho_{(\lambda,\mu)}$, and one gets the first commutative diagram in (4.3.3) below. Using Remark 3.2.3, one similarly sees that $S^{\left\{\lambda,\mu\right\}}]^{*,\operatorname{Lie}}\cong S^{\left\{\lambda^{*},\mu^{*}\right\}}$ as $\mathfrak{d}_{n}$-modules, and deduces that $\rho_{\left\{\lambda^{*},\mu^{*}\right\}}$ factors through $\rho_{\left\{\lambda,\mu\right\}}$, yielding the second commutative diagram in (4.3.3).
@@
$S^{\left\{\lambda,\pm\right\}}]^{*,\operatorname{Lie}}\cong\begin{cases}S^{\left\{\lambda^{*},\pm\right\}}&\text{if $n/2$ is even,}\\ S^{\left\{\lambda^{*},\mp\right\}}&\text{if $n/2$ is odd.}\end{cases}$
@@ 4.3.2. Self-dual Lie algebra representations
The arguments in this section are based on those in [Marin:2001, III.3.2.3]. First let $(\lambda,\mu)\in\mathcal{BP}(n)$ with $(\lambda,\mu)=(\mu^{*},\lambda^{*})$. Then $\mu=\lambda^{*}$, $n$ is even, and $\lambda$ and $\mu$ are partitions of $n/2$. Let $\langle-,-\rangle_{(\lambda,\mu)}$ be the bilinear form on $S^{(\lambda,\mu)}$ discussed in Remark 2.4.2, and let $(-|-)_{(\lambda,\mu)}$ be the nondegenerate bilinear form on $S^{(\lambda,\mu)}$ defined in terms of the linear isomorphism $\phi_{\varepsilon}=\phi_{\varepsilon}^{(\lambda,\mu)}:S^{(\lambda,\mu)}\to S^{(\lambda,\mu)}$ of Remark 2.4.5 by
@@
The bilinear form (|)(-|-) is symmetric if $n/2$ is even and is anti-symmetric if $n/2$ is odd. Thus,
@@
This implies first for $s$ in the set (2.1.1) of generators for $\mathfrak{b}_{n}$, and then for all $x\in\mathfrak{b}_{n}$ by linearity, that $(x.u|v)+(u|x.v)=0$. Next consider the basis (2.4.1) for $S^{(\lambda,\mu)}$. One has $\phi_{\varepsilon}(c_{T})=(-1)^{\ell(T^{\natural})}\cdot c_{(T^{\natural})^{*}}$, so $(c_{S}|c_{T})\neq 0$ if and only if $S=(T^{\natural})^{*}$. Let $\sigma\in\mathfrak{S}_{n}$ be the permutation that maps $T\in\mathbb{T}(\lambda,\mu)$ to $(T^{\natural})^{*}$. Then $\sigma=\sigma^{-1}$ is a product of $n/2$ disjoint transpositions, independent of $T$. Since (2.4.1) is an orthonormal basis for $S^{(\lambda,\mu)}$ with respect to the form $\langle-,$-$\rangle$, then $\sigma.c_{T}-\langle\sigma.c_{T},c_{(T^{\natural})^{*}}\rangle\cdot c_{(T^{\natural})^{*}}$ is a linear combination of vectors $c_{S}$ with $S\neq(T^{\natural})^{*}$. Further, $\langle\sigma.c_{T},c_{(T^{\natural})^{*}}\rangle\neq 0$ as a consequence of Lemma 2.4.6. Then it follows that
@@
Since $\varepsilon(\sigma)=(-1)^{n/2}$, this implies that the form (|)(-|-) is symmetric if $n/2$ is even, and is anti-symmetric if $n/2$ is odd. Finally, (|)(-|-) is non-degenerate because $\langle-,$-$\rangle$ is non-degenerate and $\phi_{\varepsilon}$ is a linear isomorphism. ∎
Next let $\left\{\lambda,\mu\right\}\in\mathcal{BP}\{n\}$ with $\left\{\lambda,\mu\right\}=\left\{\lambda^{*},\mu^{*}\right\}$. If $(\lambda,\mu)=(\mu^{*},\lambda^{*})$, then $\mathfrak{d}_{\left\{\lambda,\mu\right\}}\subseteq\mathfrak{osp}\{\lambda,\mu\}$ as defined in Lemma 4.3.1, so suppose that $(\lambda,\mu)=(\lambda^{*},\mu^{*})$. Let $(-|-)_{\left\{\lambda,\mu\right\}}$ be the nondegenerate bilinear form on $S^{\left\{\lambda,\mu\right\}}=\operatorname{Res}^{\mathcal{B}_{n}}_{\mathcal{D}_{n}}(S^{(\lambda,\mu)})$ defined via the map $\phi_{\varepsilon^{\prime\prime}}=\phi_{\varepsilon^{\prime\prime}}^{(\lambda,\mu)}$ of Lemma 2.4.4 by
@@
Let $(\lambda,\mu)\in\mathcal{BP}(n)$ with $(\lambda,\mu)=(\lambda^{*},\mu^{*})$, and let $(-|-)=(-|-)_{\left\{\lambda,\mu\right\}}$ be the bilinear form on $S^{\left\{\lambda,\mu\right\}}$ defined by (4.3.9). Then
@@
Let $b(\lambda)$ and $b(\mu)$ be the lengths of the diagonals of the Young diagrams of $\lambda$ and $\mu$, respectively. Then (|)(-|-) is symmetric if $\frac{n-b(\lambda)-b(\mu)}{2}$ is even, and is anti-symmetric otherwise. [^2] If $\lambda=\lambda^{*}$ and $\mu=\mu^{*}$, then $|\lambda|-b(\lambda)$ and $|\mu|-b(\mu)$ must each be even, hence their sum $n-b(\lambda,\mu)$ is even. Thus
@@
The proof is parallel to that of Lemma 4.3.1. In determining whether (|)(-|-) is symmetric or anti-symmetric, let $\sigma\in\mathfrak{S}_{n}\subset\mathcal{D}_{n}$ be the permutation that maps $T\in\mathbb{T}(\lambda,\mu)$ to $T^{*}$. Then $\sigma$ is a product of $\frac{|\lambda|-b(\lambda)}{2}+\frac{|\mu|-b(\mu)}{2}$ disjoint transpositions, independent of $T$, so $\varepsilon^{\prime\prime}(\sigma)=1$ if $\frac{n-b(\lambda)-b(\mu)}{2}$ is even, and $\varepsilon^{\prime\prime}(\sigma)=-1$ if $\frac{n-b(\lambda)-b(\mu)}{2}$ is odd. ∎
@@
Let $\left\{\lambda,\lambda\right\}\in\mathcal{BP}\{n\}$ with $\lambda=\lambda^{*}$ and $n/2$ even, and let $(-|-)=(-|-)_{\left\{\lambda,\pm\right\}}$ be the restriction to $S^{\left\{\lambda,\pm\right\}}$ of the bilinear form on $S^{\left\{\lambda,\lambda\right\}}$ defined by (4.3.9). Then
@@
Let $b(\lambda)$ be the length of the diagonal of the Young diagram of $\lambda$. Then (|)(-|-) is symmetric if $b(\lambda)$ is even, and is anti-symmetric if $b(\lambda)$ is odd. Thus
@@
If $\left\{\lambda,\lambda\right\}\in\mathcal{BP}\{n\}$ and $\lambda=\lambda^{*}$, then (4.3.7) and (4.3.9) define two different bilinear forms on $S^{\left\{\lambda,\lambda\right\}}=\operatorname{Res}^{\mathcal{B}_{n}}_{\mathcal{D}_{n}}(S^{(\lambda,\lambda)})$. For $n/2$ even, one can check that with respect to either form, the subspaces $S^{\left\{\lambda,+\right\}}$ and $S^{\left\{\lambda,-\right\}}$ of $S^{\left\{\lambda,\lambda\right\}}$ are orthogonal, the restrictions of the two forms to $S^{\left\{\lambda,+\right\}}$ are the same, and the restrictions of the two forms to $S^{\left\{\lambda,-\right\}}$ differ by the scalar 1-1. Thus, for $n/2$ even, the ambiguity in definition does not affect the conclusions of Lemma 4.3.3. For $n/2$ odd, one gets that with respect to either form, the spaces $S^{\left\{\lambda,+\right\}}$ and $S^{\left\{\lambda,-\right\}}$ are each totally isotropic, and the forms each define perfect pairings (differing by the scalar 1-1) between $S^{\left\{\lambda,+\right\}}$ and $S^{\left\{\lambda,-\right\}}$.